OpenStudy (anonymous):

what is the simplified form of 12x/x^2+6x+9/4x^2/x^2-9

jhonyy9 (jhonyy9):

x2 +6x +9=(x+3) squared x2 -9=(x-3)(x+3)

jhonyy9 (jhonyy9):

first fraction multiply by secondly inversed

jhonyy9 (jhonyy9):

and you can simplify (x+3) and 12x by 4x2

jhonyy9 (jhonyy9):

12x by 4x2 will remain 3/x

jhonyy9 (jhonyy9):

3/x * (x-3)/(x+3) so than

jhonyy9 (jhonyy9):

3(x-3)/x(x+3)

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

kk!

jhonyy9 (jhonyy9):

good luck bye